MATHEMATICAL ASSOCIATION OF AMERICA webwork / mat-521-spring-2018-griffin / homework-6 / 7 Homework 6: Problem 7 Previous Pr (2 points) Let X~Bin(5,0.4) Find E(7*) List Next ms You have attempted this problem 0 times You have 10 attempts remaining Page generated
Answer #1

Ans:

X has binomial distribution:

n=5,

p=0.4

E[X]=np=5*0.4=2

E[7^X]=7^2=49

E[7x]=49